# Quantifying the Balinski-Young Theorem: Structure and Probability of Quota Violations in Divisor Methods for Three States

### v2

@@ Abstract.
The apportionment problem asks how to assign representation to states based on their populations. That is, given census data and a fixed number of seats, how many seats should each state be assigned? Various algorithms exist to solve the apportionment problem, but by the Balinski-Young Theorem, every such algorithm will be flawed in some way. This paper focuses on divisor methods of apportionment, where the possible flaws are known as quota violations. This paper presents a detailed analysis of quota violations that can arise under divisor methods for three states. The study focuses on quota violations in Adams’s, Jefferson’s, Dean’s, and the Huntington-Hill methods when allocating $M$ seats. Theoretical results are proved about the behavior of these methods, particularly focusing on the types of quota violations that may occur, their frequency, and their structure. The paper then introduces tests to detect quota violations, which are then employed to construct a probability function which calculates the likelihood of such violations occurring given an initial three state population vector whose components follow varying distributions.
1. Introduction
@@
The apportionment problem - dividing a fixed number of indivisible resources among states in proportion to their populations - lies at the intersection of political science, combinatorics, and probability. A key class of apportionment methods are divisor methods, which are a group of simple methods that avoid unfair results known as apportionment paradoxes [BY01]. Divisor methods have been widely employed in the United States for assigning seats in House of Representatives, with Jefferson’s, Webster’s method, or the Huntington-Hill methods used for most of the country’s history. The Huntington-Hill method is currently used to apportion the House of Representatives.
@@
While divisor methods avoid paradoxes, these methods do not always satisfy the quota rule, which requires that each state receive an allocation equal to either the floor or ceiling of its exact proportional share. This phenomenon of imperfect apportionment is in fact guaranteed by the Balinski-Young Theorem [BY01], which establishes that no apportionment method can simultaneously satisfy the quota rule and avoid paradoxes.
@@
Since divisor methods can violate quota, and as flawed methods are unavoidable, a good question to then ask is how often will a quota violation happen, and can quota violations be predicted in advance given states’ populations? This paper answers that question for three states. Specifically, this paper investigates the frequency and structure of quota violations under divisor methods in three states in the methods of Adams’s, Jefferson’s, Webster’s, Dean’s, and Huntington-Hill. The paper characterizes how quota violations occur and provides formulas to calculate the probability that a divisor method violates quota.
2. Preliminary Notation and Definitions
@@
Given a strictly increasing divisor function $d:\mathbb{N}\to\mathbb{R}_{\geq 0}$, populations $(p_{1},\dots,p_{n})$ and a fixed number of seats $M$, an apportionment function $A(p_{1},\dots,p_{n})=(a_{1},\dots,a_{n})$, is then defined as follows:
@@
The most common and arguably reasonable divisor methods are the “five workable methods” [BY77, BY01, Hunt21] defined by the following divisor functions:
@@
The Balinski-Young Theorem [BY01] guarantees that all apportionment methods either have quota violations or have strange, unfair occurrences known as apportionment paradoxes. Divisor methods have quota violations but avouid paradoxes [BY01].
3. Results
The main results of the paper are:
@@ 3. Results
For $M$ seats and $n=3$ states, if $A$ is a divisor apportionment method with divisor function $d(s)\in[s,s+\frac{1}{2}]$ for all $s$, then upper quota violations cannot occur. In particular, upper quota violations do not occur in Adams’s, Dean’s, and Huntington-Hill’s method for $n=3$ states.
@@
(Classification of Lower Quota Violations in 3 States) Order the populations $p_{1}<p_{2}<p_{3}$. If $A$ is Adams’s, Dean’s, or the Huntington Hill method and $A(p_{1},p_{2},p_{3})$ has a quota violation, the final apportionment must be $(a_{1},a_{2},a_{3})=(\lceil q_{1}\rceil,\lceil q_{2}\rceil,\lfloor q_{3}\rfloor-1)$, where $q_{i}$ is the standard quota of state $i$. That is, the quota violation must be a lower quota violation and the lower quota violation must occur on the third state.
@@
(Lower Quota Violation Criteria Test) Let $A$ be Adams’s, Dean’s, or the Huntington-Hill method, with divisor function $d(s)$. Order the populations such that $p_{1}<p_{2}<p_{3}$ and calculate the standard quotas $q_{i}$. Then the apportionment $A(p_{1},p_{2},p_{3})$ on $M$ seats has a lower quota violation if and only if all three of the following statements are true:
  1. (1)
@@
(Classification of Upper Quota Violations in 3 States) Order the populations $p_{1}<p_{2}<p_{3}$. If $A$ is Jefferson’s method and $A(p_{1},p_{2},p_{3})$ has a quota violation, the final apportionment must be $(a_{1},a_{2},a_{3})=(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor,\lceil q_{3}\rceil+1)$, where $q_{i}$ is the standard quota of state $i$. That is, the quota violation must be an upper quota violation and the upper quota violation must occur on the third state.
@@
(Upper Quota Violation Criteria Test) Let $A$ be Jefferson’s method, with divisor function $d(s)$. Order the populations such that $p_{1}<p_{2}<p_{3}$ and calculate the standard quotas $q_{i}$. Then the apportionment $A(p_{1},p_{2},p_{3})$ on $M$ seats has an upper quota violation if and only if all three of the following statements are true:
  1. (1)
@@
  3. (3)
    $\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\lceil q_{3}\rceil=M-1$ (equivalently for $q_{1}\in\mathbb{Z}$, $\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\lfloor q_{3}\rfloor=M-2$.)
@@
(Probability of Upper Quota Violation) For $M$ seats and $n=3$ states, let $q_{1}$ and $q_{2}$ be the standard quotas of states 1 and 2 with joint pdf $f$, and define $q_{3}=M-q_{1}-q_{2}$. For $q_{1},q_{2}<q_{3}$ and Jefferson’s method, the probability of a upper quota violation is given by:
@@ 4.1. Assumptions
4.2. Introductory Lemmas and Definitions
@@
(Homogeneity) A divisor method is preserved under proportionality. If $A(p_{1},\dots,p_{n})=(a_{1},\dots,a_{n})$ then for any scalar, $\lambda\in\mathbb{R}_{>0}$, $A(\lambda p_{1},\dots,\lambda p_{n})=(a_{1},\dots,a_{n})$. In particular, apportioning the standard quotas is the same as apportioning the populations: $A(p_{1},\dots,p_{n})=A(q_{1},\dots,q_{n})$. [BY01].
@@
The equivalence of the apportionment methods $A(p_{1},\ldots,p_{n})$ in Definition 2.1 and $\tilde{A}(p_{1},\ldots,p_{n})$ is classical. See [BY01], [Hunt21], and [Pukelsheim2017].
@@
No divisor method will have both upper and lower quota violations for any $M$. [BY01, KS18].
@@
(No Violation Theorem) From [BY01], a summary of cases when no quota violations occur is as follows:
@@
    For $n=3$ states, Webster’s Method is the unique method such that upper or lower quota violations do not occur.
@@
    Adams’s Method is the unique method where upper quota violations do not occur for all $n$.
@@
    Jefferson’s Method is the unique method where lower quota violations do not occur for all $n$.
5. Structure and Probability of Quota Violations for Three States in Adams’s, Dean’s, and Huntington-Hill’s Method
We start by providing a detailed study of lower quota violations in Adams’s, Dean’s, and Huntington-Hill’s method when apportioning three states. Webster’s and Jefferson’s method have no lower quota violations by Theorem 4.5.
5.1. Characterization of Quota Violations in Three States
By the following theorem, only lower quota violations can occur in Adams’s, Dean’s, and Huntington-Hill’s method for three states.
@@ 5.1. Characterization of Quota Violations in Three States
For $M$ seats and $n=3$ states, if $A$ is a divisor apportionment method with divisor function $d(s)\in[s,s+\frac{1}{2}]$ for all $s$, then upper quota violations cannot occur. In particular, upper quota violations do not occur in Adams’s, Dean’s, and Huntington-Hill’s method for $n=3$ states.
@@
The proof above allows for lower quota violations to occur in $n=3$ states and indeed they do. The following three lemmas completely classify the lower quota violations in Adams’s, Dean’s and Huntington-Hill’s apportionment methods.
@@
In Adams’, Webster’s, and Huntington-Hill’s methods for $n=3$ states, lower quota violations can only occur on the state with the largest population.
@@
Let $d$ be the divisor function of Adams’s, Dean’s, or Huntington-Hill’s method. Then for $a,b\in[1,\infty)$, $a\cdot d(\lfloor b\rfloor)<b\cdot d(\lfloor a\rfloor-1)$ implies $a\geq b$.
@@
For Dean’s method, let $d(s)=\frac{2s(s+1)}{2s+1}$. First, notice if $a<2$, then $a\cdot d(\lfloor b\rfloor)<b\cdot d(\lfloor a\rfloor)$ never holds. Thus $a\geq 2$. Additionally, this implies the statement is true if $b<2$. The proof for $a,b\geq 2$ will be done by showing the contrapositive. Let $2\leq a\leq b$. Let $n=\lfloor a\rfloor,m=\lfloor b\rfloor$. Note $m\geq n\geq 1$. Define $f(x)=\dfrac{x}{d(\lfloor x\rfloor-1)}$ and $g(x)=\dfrac{x}{d(\lfloor x\rfloor)}$. Then, it suffices to show $f(a)\geq g(b)$.
@@
For the Huntington-Hill method or Adams’s method, the proof is similar.
@@
(Classification of Lower Quota Violations in 3 States) Order the populations $p_{1}<p_{2}<p_{3}$. If $A$ is Adams’s, Dean’s, or Huntington-Hill’s method and $A(p_{1},p_{2},p_{3})$ has a quota violation, the final apportionment must be $(a_{1},a_{2},a_{3})=(\lceil q_{1}\rceil,\lceil q_{2}\rceil,\lfloor q_{3}\rfloor-1)$, where $q_{i}$ is the standard quota of state $i$. That is, the quota violation must be a lower quota violation and the lower quota violation must occur on the third state.
5.2. Lower Quota Violation Criteria Test for 3 States
@@
(Lower Quota Violation Criteria Test) Let $A$ be Adams’s, Dean’s, or the Huntington-Hill method, with divisor function $d(s)$. Order the populations such that $p_{1}<p_{2}<p_{3}$ and calculate the standard quotas $q_{i}$. Then the apportionment $A(p_{1},p_{2},p_{3})$ on $M$ seats has a lower quota violation if and only if all three of the following statements are true:
  1. (1)
@@
5.3. Probability Function for Lower Quota Violations in Three States
For a fixed $M$, $n=3$ states, and with $A$ Adams’s, Dean’s, or the Huntington-Hill method, this section uses the Lower Quota Violation Criteria Test (5.6) to find the probability of a lower quota violation.
@@ 5.3. Probability Function for Lower Quota Violations in Three States
5.3.1. Motivating Example
As an example, let $M=100,\lfloor q_{1}\rfloor=1,\lfloor q_{2}\rfloor=2$, and $A$ the Huntington-Hill method. Write $q_{1}=\lfloor q_{1}\rfloor+d_{1}$ and $q_{2}=\lfloor q_{2}\rfloor+d_{2}$, where $d_{1},d_{2}$ are the decimal parts. The feasible region of the Lower Quota Violation Criterion Test is graphed below;

### v3

@@ Abstract
The apportionment problem asks how to assign representation to states based on their populations. That is, given census data and a fixed number of seats, how many seats should each state be assigned? Various algorithms exist to solve the apportionment problem, but by the Balinski-Young Impossibility Theorem, every such algorithm will be flawed in some way. This paper focuses on divisor methods of apportionment, where the possible flaws are known as quota violations.
This paper presents a detailed analysis of quota violations that can arise under divisor methods for three states, By focusing on the three-state case, the paper makes the consequences of the Balinski–Young theorem particularly transparent and allows for a precise classification of quota violations that is difficult to obtain in more general formulations. The study focuses on quota violations in the Adams, Jefferson, Dean, and the Huntington-Hill methods when allocating $M$ seats, but is expandable to a wider class of divisor functions. Theoretical results are proved about the behavior of these methods, particularly focusing on the types of quota violations that may occur, their frequency, and their structure and geometry. The key results of the paper are tests to detect quota violations which are employed to understand the geometry of violations and construct a probability function which calculates the likelihood of such violations occurring given an initial three state population vector whose components follow varying distributions.
1 Introduction
@@
The apportionment problem - dividing a fixed number of indivisible resources among states in proportion to their populations - lies at the intersection of political science, combinatorics, and probability. A key class of apportionment methods are divisor methods, which are a group of simple methods that avoid unfair results known as apportionment paradoxes [undefa]. Divisor methods have been widely employed in the United States for assigning seats in House of Representatives, with the Jefferson, Webster, or Huntington-Hill method used for most of the country’s history. The Huntington-Hill method is currently used to apportion the House of Representatives. [undefe]
@@
While divisor methods avoid paradoxes, these methods do not always satisfy the quota rule, which requires that each state receive an allocation equal to either the floor or ceiling of its exact proportional share. This phenomenon of imperfect apportionment is in fact guaranteed by the Balinski-Young Impossibility Theorem [undefa], which establishes that no apportionment method can simultaneously satisfy the quota rule and avoid paradoxes.
@@
Since divisor methods can violate quota, and as flawed methods are unavoidable, good questions to ask are: What is the underlying structure of such violations? How often will a quota violation happen? And, can quota violations be predicted in advance given states’ populations? This paper answers these question for three states. Specifically, The paper investigates the frequency, structure, and geometry of quota violations under divisor methods in three states in the Adams, Jefferson, Webster, Dean, and Huntington-Hill methods. The paper characterizes how quota violations occur and provides formulas to calculate the probability that a divisor method violates quota. Overall, this paper provides a detailed analysis of quota violations in the simpler case of three states, which ultimately highlights and corroborates the Balinski Young Impossibility Theorem.
This paper builds on related work about probabilities and analysis of apportionment within social choice theory and the mathematics of elections, such as similar probabilistic work on simplexes concerning the Jagiellonian compromise found in [undefo] and [undefn]; work on systemic bias such as in [undefl], [undefm], [undefb], and [undefh]; and work on the probabilities of voting paradoxes such as in [undefc] and [undefj].
2 Preliminary Notation and Definitions
@@
(Divisor Method) Given a strictly increasing divisor function $d:\mathbb{Z}_{\geq 0}\to\mathbb{R}_{\geq 0}$, populations $(p_{1},\dots,p_{n})$ and a fixed number of seats $M$, an apportionment function $A(p_{1},\dots,p_{n})=(a_{1},\dots,a_{n})$, is then defined as follows:
@@
The most common and arguably reasonable divisor methods are the “five workable methods” [undef, undefa, undefg] defined by the following divisor functions:
@@
The Balinski-Young Theorem [undefa] guarantees that all apportionment methods either have quota violations or have strange, unfair occurrences known as apportionment paradoxes. Divisor methods have quota violations but avoid paradoxes [undefa].
3 Results
The following is a summary of the results of the paper:
@@ 3 Results
For $M$ seats and $n=3$ states, if $A$ is a divisor apportionment method with divisor function $d(s)\in[s,s+\frac{1}{2}]$ for all $s$, then upper quota violations cannot occur. In particular, upper quota violations do not occur in the Adams, Dean, and Huntington-Hill methods for $n=3$ states.
@@
(Classification of Lower Quota Violations in 3 States) Order the populations $p_{1}<p_{2}<p_{3}$. If $A$ is the Adams, Dean, or the Huntington Hill method and $A(p_{1},p_{2},p_{3})$ has a quota violation, the final apportionment must be $(a_{1},a_{2},a_{3})=(\lceil q_{1}\rceil,\lceil q_{2}\rceil,\lfloor q_{3}\rfloor-1)$, where $q_{i}$ is the standard quota of state $i$. That is, the quota violation must be a lower quota violation and the lower quota violation must occur on the third state.
@@
(Lower Quota Violation Criteria Test) Let $A$ be the Adams, Dean, or the Huntington-Hill method, with divisor function $d(s)$. Order the populations such that $p_{1}<p_{2}<p_{3}$ and calculate the standard quotas $q_{i}$. Then the apportionment $A(p_{1},p_{2},p_{3})$ on $M$ seats has a lower quota violation if and only if all three of the following statements are true:
  1. 1.
@@
(Classification of Upper Quota Violations in 3 States) Order the populations $p_{1}<p_{2}<p_{3}$. If $A$ is the Jefferson method and $A(p_{1},p_{2},p_{3})$ has a quota violation, the final apportionment must be $(a_{1},a_{2},a_{3})=(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor,\lceil q_{3}\rceil+1)$, where $q_{i}$ is the standard quota of state $i$. That is, the quota violation must be an upper quota violation and the upper quota violation must occur on the third state.
@@
(Upper Quota Violation Criteria Test) Let $A$ be the Jefferson method, with divisor function $d(s)$. Order the populations such that $p_{1}<p_{2}<p_{3}$ and calculate the standard quotas $q_{i}$. Then the apportionment $A(p_{1},p_{2},p_{3})$ on $M$ seats has an upper quota violation if and only if all three of the following statements are true:
  1. 1.
@@
  3. 3.
    $\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\lceil q_{3}\rceil=M-1$ (equivalently for $q_{3}\notin\mathbb{Z}$, $\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\lfloor q_{3}\rfloor=M-2$.)
@@
(Probability of Upper Quota Violation) For $M$ seats and $n=3$ states, let $q_{1}$ and $q_{2}$ be the standard quotas of states 1 and 2 with joint pdf $f$, and define $q_{3}=M-q_{1}-q_{2}$. For $q_{1},q_{2}<q_{3}$ and the Jefferson method, the probability of a upper quota violation is given by:
@@ 4.1 Assumptions
4.2 Introductory Lemmas and Definitions
@@
[undefa] (Homogeneity) A divisor method is preserved under proportionality. If $A(p_{1},\dots,p_{n})=(a_{1},\dots,a_{n})$ then for any scalar, $\lambda\in\mathbb{R}_{>0}$, $A(\lambda p_{1},\dots,\lambda p_{n})=(a_{1},\dots,a_{n})$. In particular, apportioning the standard quotas is the same as apportioning the populations: $A(p_{1},\dots,p_{n})=A(q_{1},\dots,q_{n})$. .
@@
The equivalence of the apportionment methods $A(p_{1},\ldots,p_{n})$ in Definition 2.1 and $\tilde{A}(p_{1},\ldots,p_{n})$ is classical. See [undefa], [undefg], and [undefk].
@@
[undefa, undefi] No divisor method will have both upper and lower quota violations for any $M$.
@@
[undefa] (No Violation Theorem) From , a summary of cases when no quota violations occur is as follows:
@@
    For $n=3$ states, the Webster method is the unique method such that upper or lower quota violations do not occur.
@@
    The Adams Method is the unique method where upper quota violations do not occur for all $n$.
@@
    The Jefferson Method is the unique method where lower quota violations do not occur for all $n$.
5 Structure and Probability of Quota Violations for Three States in the Adams, Dean, and Huntington-Hill Methods
We start by providing a detailed study of lower quota violations in the Adams, Dean, and Huntington-Hill methods when apportioning three states. The Webster and Jefferson methods have no lower quota violations by Theorem 4.5.
5.1 Characterization of Quota Violations in Three States
By the following theorem, only lower quota violations can occur in the Adams, Dean, and Huntington-Hill methods for three states.
@@ 5.1 Characterization of Quota Violations in Three States
For $M$ seats and $n=3$ states, if $A$ is a divisor apportionment method with divisor function $d(s)\in[s,s+\frac{1}{2}]$ for all $s$, then upper quota violations cannot occur. In particular, upper quota violations do not occur in the Adams, Dean, and Huntington-Hill methods for $n=3$ states.
@@
The proof above allows for lower quota violations to occur in $n=3$ states and indeed they do. The following three lemmas completely classify the lower quota violations in the Adams, Dean and Huntington-Hill apportionment methods.
@@
In the Adams, Webster, and Huntington-Hill methods for $n=3$ states, lower quota violations can only occur on the state with the largest population.
@@
Let $d$ be the divisor function of the Adams, Dean, or Huntington-Hill method. Then for $a,b\in[1,\infty)$, $a\cdot d(\lfloor b\rfloor)<b\cdot d(\lfloor a\rfloor-1)$ implies $a\geq b$.
@@
For the Dean method, let $d(s)=\frac{2s(s+1)}{2s+1}$. First, notice if $a<2$, then $a\cdot d(\lfloor b\rfloor)<b\cdot d(\lfloor a\rfloor)$ never holds. Thus $a\geq 2$. Additionally, this implies the statement is true if $b<2$. The proof for $a,b\geq 2$ will be done by showing the contrapositive. Let $2\leq a\leq b$. Let $n=\lfloor a\rfloor,m=\lfloor b\rfloor$. Note $m\geq n\geq 1$. Define $f(x)=\dfrac{x}{d(\lfloor x\rfloor-1)}$ and $g(x)=\dfrac{x}{d(\lfloor x\rfloor)}$. Then, it suffices to show $f(a)\geq g(b)$.
@@
For the Huntington-Hill method or the Adams method, the proof is similar.
@@
(Classification of Lower Quota Violations in 3 States) Order the populations $p_{1}<p_{2}<p_{3}$. If $A$ is the Adams, Dean, or Huntington-Hill method and $A(p_{1},p_{2},p_{3})$ has a quota violation, the final apportionment must be $(a_{1},a_{2},a_{3})=(\lceil q_{1}\rceil,\lceil q_{2}\rceil,\lfloor q_{3}\rfloor-1)$, where $q_{i}$ is the standard quota of state $i$. That is, the quota violation must be a lower quota violation and the lower quota violation must occur on the third state.
5.2 Lower Quota Violation Criteria Test for 3 States
@@
(Lower Quota Violation Criteria Test) Let $A$ be the Adams, Dean, or Huntington-Hill method, with divisor function $d(s)$. Order the populations such that $p_{1}<p_{2}<p_{3}$ and calculate the standard quotas $q_{i}$. Then the apportionment $A(p_{1},p_{2},p_{3})$ on $M$ seats has a lower quota violation if and only if all three of the following statements are true:
  1. 1.
@@
Remark 5.7 (Generalization to $n$ states).
Although Theorem 5.6 is stated for three states, the underlying ideas extend in principle to $n\geq 4$. The main complication already appears at $n=4$: there is no analogue of Theorem 5.5 providing a single canonical form for quota violations. Instead, one must account for multiple distinct configurations, each requiring its own criterion, leading to a disjunction of conditions. The number of such cases grows rapidly with $n$, making the resulting criterion unwieldy. For clarity and focus, we restrict attention to the case $n=3$ and leave the general $n$-state setting for future work.
5.3 Geometric Visualization of Lower Quota Violations
Using the Lower Quota Violation Criteria Test (5.6) and Theorem 5.8, we can display all quota violations on the simplex of standard quotas $\{(q_{1},q_{2},q_{3})\mid q_{1}+q_{2}+q_{3}=M\}$ within the Adams, Dean, or Huntington-Hill method. This creates a visualization of the geometry and structure of lower quota violations. For instance, below is a simplex plot of all violatory $(q_{1},q_{2},q_{3})$ in the Huntington-Hill method with $M=13$:
Figure 1: Quota Violations in the Huntington-Hill Method on $\{(q_{1},q_{2},q_{3})\mid q_{1}+q_{2}+q_{3}=M\}$
5.4 Probability Function for Lower Quota Violations in Three States
For a fixed $M$, $n=3$ states, and with $A$ the Adams, Dean, or Huntington-Hill method, this section uses the Lower Quota Violation Criteria Test (5.6) to find the probability of a lower quota violation.
@@ 5.4 Probability Function for Lower Quota Violations in Three States
5.4.1 Motivating Example
As an example, let $M=100,\lfloor q_{1}\rfloor=1,\lfloor q_{2}\rfloor=2$, and $A$ the Huntington-Hill method. Write $q_{1}=\lfloor q_{1}\rfloor+d_{1}$ and $q_{2}=\lfloor q_{2}\rfloor+d_{2}$, where $d_{1},d_{2}$ are the decimal parts. The feasible region of the Lower Quota Violation Criterion Test is graphed below:
Figure 2: Example Inequality Region with Overlap